I need to help my daughter with her algebra and show my work.
I know x=7 but can't help her show how I got there
5(x+2)=6x +3

Answer by ewatrrr(24785)


 
Hi,
5(x+2)=6x +3 |Treat Parenthesis 1st
5x + 10 = 6x + 3
Isolating variable on one side of the EQ
5x -5x+ 10 = 6x-5x + 3 |Subtracting 5x from both sides of the EQ
10 = x + 3
10-3 = x +3 - 3 | Subtracting 3 from both sides of the EQ
7 = x